Please help me solve this problem.




A card is selected at random from a deck of 52 cards. What are the odds in favor of the card being a jack? State answer in ratio format (1 to 1)





Answer by sudhanshu_kmr(1152)


no. of jack = 4
probability of card being jack P(A) = 4/52 = 1/13
P'(A) = 12/13
odds in favor = P(A) / P'(A) = 1/12